HUBBLE SPACE TELESCOPE DAILY REPORT #1053 PERIOD COVERED: 0700 EST Daily Status Report as of 014/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 4721 (WF/PC-2 Decontamination and TEC Turn On) The WF/PC-2 was used to make ultraviolet photometric observations of the standard star A+81D266, plus K-Spots, and dark calibration images. All four sets of CCDs were used. A total of 11 images were received and displayed in the OSS. No anomalies were reported. 1.2 Completed WF/PC-2 4763 (WF/PC-2 QE and Throughput Measurement) The WF/PC-2 was used to make ultraviolet exposures of the UV flux-standard star BD+75D325 to be used for alignment studies. The OSS monitored the observations, and assessed the data quality. No anomalies were noted. 1.3 Completed Two Sets of FOS/RD/BL 4746 (COSTAR/FOS Coarse Alignment) The FOS mirror adjustments were successfully computed and uplinked to the spacecraft. The Faint Object spectrograph was used to make pairs of acquisitions, one from each side (i.e., Red and Blue), to be used for coarse alignment and focus verification of the COSTAR FOS M1 mirror. The cluster NGC-188-225 was used as a target object. No problems were reported. 1.4 Completed WF/PC-2 5502 (ERO Imaging of QSO Host Galaxies) The WF/PC-2 was used to make images at two different continuum wavelengths of the bright inner region of the low redshift quasi-stellar source 1229+204. The baseline guide star acquisition got fine lock. A total of eight images were received and displayed. The observations were reported to look nominal. 1.5 Completed Two Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Field Calibration) The WF/PC-2 was used to make a set of visual internal flat images. Four images were received and displayed. One of the images was reported to be underexposed, and the others were nominal. 1.6 Completed FOC/96 4787 (Observations of an External Seyfert Galaxy) The Faint Object Camera (f/69) and COSTAR were used to make observations of standard star BPM6274 to help measure the system's throughput. The baseline acquisition got fine lock. A corrective SIP maneuver was successfully uplinked to the spacecraft. All observations were reported to be nominal. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled :11 Successful :11 2.3 Significant HSTARs: None 2.4 Operations Notes: All HST operations, whether for SMOV activities or for normal health and safety monitoring, continue to be nominally performed. Further COSTAR positioning for FOS operations was accomplished. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 4385 (Spectroscopy and Imagery of an External Source; Cycle 3 High) 3.2 Complete Four Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Fields Calibration) 3.3 Complete Three Sets of FOS/RD/BL 4746 (COSTAR/FOS Coarse Alignment) 3.4 Complete WF/PC-2 4616 (WF/PC-2 Fine Alignment and Focus) 3.5 Complete GHRS/2 5198 (The Evolution of Boron: Implications for Galactic Evolution) 3.6 Complete FGS 5054 (High Speed Astrometry - A Search for Planetary Companions to Low Mass Stars)) 3.7 Complete Three Sets of WF/PC-2 4721 (WF/PC-2 Decontamination and TEC Turn ON) 3.8 Complete Three Sets of S/C 5254 (GHRS SMOV Side One Activation) 3.9 Complete WF/PC-2 4709 (WF/PC-2 Internal Monitor) 3.10 Complete FGS 2565 (Linking HIPPARCOS to the Extragalactic Reference Frame; Part 1 of 6) 3.11 Complete GHRS/2 3936 (Absorption Cloud Physics in Seyfert Nuclei; Cycle 2) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: A series of engineering tests will be initiated during the weekend as part of the GHRS Side One turn-on. /DMH -END-